Current Conditions

25/2/2018

The water is warm and clear. Despite it getting cold out of the water, in the water is great.

23/7/2017

The big swells made it quite green in the nursery but on Sunday the vis varied from 5 to 15m at the docks.

7/5/2017

20m+ visibility on the wreck of the Wandra. 10-15m vis at The Drum wit the seals, who were quite playful.

17/4/2017

A bit green in the nursery, but still 8-10m vis and lots of giant grey leather jacket action, wobblies port jacksons, rays and the usual suspects. Not so clear up at the drum, but more giant grey leather jackets, and the seals have been playful. More giant grey leatherjackets on The Wall, North Bowen.

9/4/2017

Around the docks on Sunday, it was green. We didn't get into the water. At the nursery it was about 5-8m vis. The divers were quite happy though with probably a dozen giant grey leatherjackets.

2/4/2017

After the dirty water we've been experiencing, it's finally starting to clear up. 15-18m at The Drum, 25m plus at Steamers if Warren is to be believed ;-) and 8-10m around Bowen Island. Seedy Wee Dragons, Grey Nurse and happy divers.

15/1/2017

Visibility not as good as last weekend, but still up to 10m vis around Bowen Island.

8/1/2017

10-15m vis in the nursery, 20m vis at the docks and nearly blue.

28/12/2016

Conditions a bit weird on the North Side today. In the morning the vis at Deco Rock at the Docks was about 15m, while the area around Slot Cave was 3m, this afternoon, Deco Rock was a green 4m vis, while the area around Slot Cave was up to 15m vis and not green. Pretty interesting for two sites within a few hundred metres of each other. 18-20 degrees.

18/12/2016

Up to 10m vis in the nursery, 15m vis on the wall at North Bowen. About 18 degrees C.

11/12/2016

Green and snotty. Unless you have some cool location that is probably deep dark and scary, you're going to be dealing with green. Inside Bowen and North of Point Perp on the outside were pretty much the same, below 30m at the Arch was ok, but we didn't have digging equipment in the nursery.

3/12/2016

15m vis and 21 degrees in the Honeymoon Bay area, even clearer on Sunday the 4th.

20/11/2016

5-12m vis on the North side, though, around Deco rocks at The Docks it was greener and less, once they got to the Double Decker Cave it cleared up a little.

6/11/2016

A very green 3-5m vis at Bowen Island, even with the incoming tide. 10-12m vis inside Point Perpendicular

28/8/2016

10-15 vis at the drum. Seals not so playful but perseverance pays off. 10-15m blue water at Ronnie's Cathedral/Labyrinth/Whorehouse.

6/8/2016

4m vis at the drum, 15m North Bowen and about 8m in the nursery

24/7/2016

12-15m vis around Spider cave, 10-12m vis at the Drum

17/7/2016

South of Bowen Island, up to 10m vis. Still lots of PJ's.

10/7/2016

Around Bowen Island it was green and snotty. Still, lots of Port Jackson sharks and in the nursery one of the divers took some nice photos of a Giant Cuttlefish. The nursery was ok but North Bowen was cold. Waz came out shivering.

3/7/2016

North of Point Perpendicular the water is still green, averaging around 5-8m vis, 16 degrees..

Around Bowen Island and to the South, 15-20m vis, 18 degrees.

17/6/2016

After the long weekend's flooding, associated runoff and big North East Swells, visibility dropped off considerably. While you could see the hand in front of your face, it was far from ideal.

Fortunately conditions are starting to return to how they were before. It's not 20 degrees and 20m vis, as it had been, however last weekend at Bowen Island we had 12-15m vis, with lots of life including mating Port Jackson sharks, big cuttlefish in abundance, schooling fishies and 16 degree water.

If you're up for a trip to the Drum, the seals are more playful than ever. The visibility hasn't been as good with patches of yuck messing things up, but if you're after a bit of an adventure, it's protected from the Southerly swells we have been getting and is clearing up as well.

If you have a trip report or update on the conditions, feel free to update our facebook page and I'll include it here.

current_conditions.txt · Last modified: 2018/03/01 08:33 by andrew